DIRC Operations Weekly Report

Period: Wednesday 2007/5/03 - Friday 2007/5/18

Oncall: Alejandro Perez
Fast Monitoring: Alejandro Perez, Jose Benitez

File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20070515.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20070515

Contents


BaBar-PEP news

PEP peak lumi: ~8.0 x 1033 (three weeks ago: ~7.0 x 1033)

BaBar integrated lumi: 0.0 fb-1 (three weeks ago: 1.8 fb-1)

Run 6 PEP delivery: 0.0 fb-1 (three weeks ago: 1.8 fb-1)

Run 6 total delivery: ~32.4 fb-1 Run 6 total BaBar recorded: 31.2 fb-1 including 0 fb-1 off-peak data

PEP activities this week:

  • 5/1 : Q4 chamber replacement started.
    Problem: Q2 failed to ramp down, Q2 melted and water hose broke when the cooling was turned off.
    Consequences: All systems powered off. Much water spilled in DIRC tunnel, flowed mostly backward but also forward.
  • 5/3: Q4 chamber taken out.
  • 5/7: After many tests on the Q2 magnet (since 5/1), PEP decided to replace it. The Q2 magnet went out.
  • 5/8-10: Characterization measurements for the new Q2 magnet .
  • 5/10: Q2 magnet and Q4 chamber succefully reinstalled.
  • 5/14: Vacuum check completed. Magnets ramped up, started injecting.
  • 5/15-16: PEP coasting beems and taking a few test runs.

    BaBar activities/issues this week:

  • Two weeks ROD starting on May 1.
  • 5/1:
  • 5/3: EMC backward electronics, DCH and DIRC turned back on succesfully.
  • 5/4:
  • 5/9:
  • 5/11: EMC replacement of a couple ADB boards.
  • 5/15: SVT turned back on.

    DIRC activities